OpenStudy (anonymous):

Help! i give medals:) The volumes of two cylinders are 2197πcubed and 8000πcubed find the following ratios:I will post them up give me a sec

OpenStudy (anonymous):

 \[\frac{ r1 }{ r2 }=\frac{ h1 }{ h2 }=\frac{ a }{ b }=\] \[\frac{ Area I }{ Area II }=\frac{ a ^{2}}{ b ^{2} }=\]